8. Apply the Picard iteration scheme to the linear system * = Ax, x(0) = Xo and show that the solution is given by x(t) = eAl pls with good handeriting
Answer #1

Consider, dx = Alt)X(+) X(to) =_Xo bulin from to to t - ☺ Integrate X(t) - X(ta) = ACO X Colds =) X(t) = x(7) + JAOX (D) do 1where $(t, to) = 1+ Amolds + .. which is olt, to) called is the called state tramition matrix Peano - Baker sesienA is a constant matsin , then If $(t, to) = I + A + A I touto do, o tama = I + Alt-to) + A Alt-to)to = 0 = $(t,0) = At = [X(t) = eAt Xo/